OpenStudy (anonymous):

cos 3x over cos x=1-4sin^2 x

OpenStudy (anonymous):

use ,\[\cos 3x=4\cos ^{3 }x-3\cos x \]

OpenStudy (anonymous):

^???????

OpenStudy (anonymous):

\[\frac{ \cos 3x }{ \cos x }=\frac{ 4\cos ^{3}x-3\cos x }{ \cos x }\]

OpenStudy (anonymous):

take cosx common from the numerator and cancel with denominator.

OpenStudy (anonymous):

@surjithayer which rule did you use to get 4cos^3x-3cosx ?

OpenStudy (anonymous):

\[then use \sin ^{2}x+\cos ^{2}x=1\]

OpenStudy (anonymous):

it is also a formula

OpenStudy (anonymous):

I'm not understanding anything

OpenStudy (anonymous):

Sketch of my solution: Let cos 3x=cos(x+2x) use addition rule, simplify. You will have terms with cos 2x and sin 2x; use the double angle rule, continue to simplify. Eventually, get \[\cos^2x+\sin^2x=1\] Done.

OpenStudy (anonymous):

\[\cos 3x=\cos \left( 2x+x \right)=\cos 2x \cos x-\sin 2x \sin x\] \[=\left( \cos ^{2}x-\sin ^{2}x \right)\cos x-2\sin x \cos x \sin x\] \[=\left\{ \cos ^{2}x-\left( 1-\cos ^{2}x \right) \right\}\cos x-2\cos x \sin ^{2}x\] \[=(\cos ^{2}x-1+\cos ^{2}x)\cos x-2\cos x \left( 1-\cos ^{2}x \right)\] \[=2\cos ^{3}x-\cos x-2\cos x+2\cos ^{3}x=4\cos ^{3}x-3\cos x\]

OpenStudy (anonymous):

i think now you know the formula for cos 3x

OpenStudy (anonymous):

What happens to 1-4sin^2 x

OpenStudy (anonymous):

\[\frac{ \cos 3x }{ \cos x }=\frac{ 4\cos ^{3}x-3\cos x }{ \cos x }=\frac{ \cos x \left( 4\cos ^{2}x-3 \right) }{\cos x }\] \[=4\cos ^{2}x-3=4\left( 1-\sin ^{2}x \right)-3=4-4\sin ^{2}x-3=1-4\sin ^{2}x\]

OpenStudy (anonymous):

\[\frac{\cos3x}{cosx}=1-4\sin^2x \implies \frac{\cos(2x+x)}{cosx}=1-4\sin^2x\]\[=\frac{\cos2xcosx-\sin2xsinx}{cosx}=\cos2x-\sin2xtanx=1-4\sin^2x\]

OpenStudy (anonymous):

\[\cos^2x-\sin^2x-2sinxcosxtanx=\cos^2x-\sin^2x-2\sin^2x=1-4\sin^2x\]\[\implies \cos^2x+\sin^2x=1~~~QED\]

OpenStudy (anonymous):

Work only on the LHS until the last step (combining like terms). It's a little tedious, but not horrible if you use your rules and pay attention to the algebra.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

No sweat. Do math every day.